my math question is like this
1/2[x -1/6]+2/3=5/6+1/3[1/2-3X]
I WANT MY ANSWERED SOON PLZ

Answer by rothauserc(4718)

1/2[x -1/6]+2/3=5/6+1/3[1/2-3X]
LCD is 6, multiply both sides of = by 6
3[x -1/6] +4 = 5 +2[1/2-3X]
3x -1/2 +4 = 5 +1 -6x
3x +7/2 = 6 -6x
9x = 5/2
x = 5/18